OpenStudy (anonymous):

What are the pair of functions representing f(x) and g(x) if f(g(x))= 4x+11? a.) f(x)=4x+7 and g(x)=x-3 b.) f(x)=4x-7 and g(x)=x+3 c.) f(x)=x-3 and g(x)=4x+7 d.) f(x)=4x+3 and g(x) x+2 Thank you so much for any help!

hartnn (hartnn):

Lets take first option: \(f(x) = 4x+7 \\ g(x) = x-3 \) to find \(f(g(x))\), just replace 'x' in f(x) by g(x) \(f(g(x)) = 4(x-3)+7 = 4x-12+7 = 4x-5\) This is not 4x+11 :) So fist option is not your answer. Can you try this for other options?

OpenStudy (anonymous):

Yay! I think I understand. Is it d? 4(x+2)+3=4x+8+3=4x+11

hartnn (hartnn):

YES. you're correct :) Good work!

OpenStudy (anonymous):

Awesome, thank you so much!

hartnn (hartnn):

most welcome ^_^ you seem to be new here. \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)